Jim's car will travel 240 mi. on 15 gal. of gas. How far can he expect to go on 3 gal. of gas?
Any ideas? Thanks so much.

Answer by checkley77(12844)

240/15=X/3
15X=3*240
15X=720
X=720/15
X=48 MILES CAN BE DRIVEN ON 3 GALLONS OF GAS.